OpenStudy (luckycoins888):

what is 1 5/8 + 3 7/9

OpenStudy (anonymous):

hmmmm

OpenStudy (luckycoins888):

hard

OpenStudy (anonymous):

well, first you have to find the greatest common factor. Do you know what that is?

OpenStudy (luckycoins888):

no

OpenStudy (anonymous):

you find the number that both 9 and 8 are a factor of which is 72. 8 goes into 72 9 times so you would multiply the numerator (the number on top) by 9 and you get 1(45/72). 9 goes into 72 8 times so would multiply that numerator by 8 and get 3(56/72). THEN you take the number in front of the fraction and say well, this is 1. So 1 times 72 is 72. Then add that number to your (45/72) and you get (117/72). The next number has a three in front of it so you do 3 time 72 which is 216 then add to (56/72) and get (272/72). SO... Now the equation is \[\frac{ 117 }{ 72 }+\frac{ 272 }{ 72 }\] now add the top two numbers and you get \[\frac{ 389 }{ 72 }\] FINALLY...simplify if you can. Does that make any sense?

OpenStudy (luckycoins888):

yes thx alot